I'm trying to work a problem on how long it will take a savings acct bal to reach 1,000,000 if i deposited 2947 at 9% interest compounded quarterly,
I'm using the equation A=P(1+i)^n
with no luck,,please help

Answer by lwsshak3(11628)

how long it will take a savings acct bal to reach 1,000,000 if i deposited 2947 at 9% interest compounded quarterly
***
A=P(1+i)^n
P=initial investment=2947
i=interest per period(quarter)=.09/4=0.0225
n=number of periods
A=amount after n periods=10^6
..
10^6=2497(1+.0225)^n
10^6/2497=(1.0225)^n
take log of both sides
log(10^6)-log(2497)=nlog(1.0225)
log(10^6)=6
6-log(2497)/(log(1.0225))=n
n≈269 periods(quarters) or ≈ 67 yrs